OpenStudy (anonymous):

If volume of a sphere is increased by 72.8% then by what % its surface area will increase?

OpenStudy (anonymous):

Help me...

OpenStudy (anonymous):

consider sa = v r^2 = r^3 if v is increased by 72.8, then R= cube root under 1.728r so sa= R^2 = (cube root under 1.728r)^2 = 1.44r thus surface area will increase by 44%

OpenStudy (anonymous):

saruz ! if sa = v then how r^2 = r^3 because sa = 4 pie r^2 and v = 4/3 pie r^3

OpenStudy (anonymous):

saruz reply soon.....

OpenStudy (anonymous):

that is not true at all ........that is the thing you need to do in rough.....its just a fast method of doing mathematics..... sa never = v but sa (three equal to sign) v

OpenStudy (anonymous):

that is an assumption!

OpenStudy (anonymous):

did i get the answer wrong?>

OpenStudy (anonymous):

your ans. is right but after assumption when we equal sa and v then (see fig.)